OpenStudy (anonymous):

What is the general form of the equation of a circle with a center at the (–10, 5) and a radius of 11? (x – 10)^2 + (y – 5)^2 = 121 (x – 10)^2 + (y + 5)^2 = 121 (x + 10)^2 + (y – 5)^2 = 121 (x + 10)^2 + (y + 5)^2 = 121

OpenStudy (anonymous):

Standard form for circle: (x-h)^2+(y-k)^2=r^2 You need to change the sign of the center(-10,5). You will get (x-(-10))^2+(y-5)^2=11^2 =(x+10)^2+(y-5)^2=121 C. is correct